# Flex3D: Feed-Forward 3D Generation With Flexible Reconstruction Model And Input View Curation

### v2

@@ 4.3 Ablation study and analysis
Here we focus on the candidate view generation and selection pipeline, utilizing a fully trained FlexRM, fine-tuned with simulated imperfect data, as the reconstruction model. The evaluation protocol follows that outlined in section 4.2. We conduct ablation experiments by removing: (1) multi-view generation at varying elevations, (2) consistency verification (resulting in random view selection), and (3) back view generation quality assessment.
Table 4 summarizes the results, demonstrating that the removal of any of these components leads to a decrease in both CLIP (Radford et al., 2021) and VideoCLIP (Wang et al., 2023) text similarity scores. This shows the contribution of each component in achieving high-quality 3D generation.
@@
We analyze our imperfect data simulation strategy using both reconstruction and generation metrics. Evaluation settings mirror those used in the ablation study. As shown in table 5, incorporating imperfect data simulation yields improvements across both generative and reconstruction tasks. This suggests that our strategy effectively exposes the model to a wider range of data variations, enhancing its overall performance and robustness.
@@ 5 Conclusion
This paper introduces Flex3D, a novel feed-forward 3D generation pipeline capable of creating high-quality 3D Gaussian representations from either text or a single image. Its core component is FlexRM, a robust and flexible reconstruction model designed to handle an arbitrary number of input views. Extensive evaluations on benchmark tasks demonstrate that Flex3D achieves state-of-the-art performance in both 3D reconstruction and generation tasks. These results highlight the effectiveness of our proposed approach in addressing the challenges of feed-forward 3D generation, paving the way to more robust and versatile 3D content creation from diverse input sources.
Ethics Statement
@@
Appendix A Implementation details

### v3

@@ 4.3 Ablation study and analysis
Here we study the candidate view generation and selection pipeline, utilizing a fully trained FlexRM, fine-tuned with simulated imperfect data, as the reconstructor. The evaluation protocol follows that outlined in section 4.2. We conduct ablation experiments by removing: (1) multi-view generation at varying elevations, (2) consistency verification (resulting in random view selection), and (3) back view generation quality assessment.
Table 4 summarizes the results, demonstrating that the removal of any of these components leads to a decrease in both CLIP (Radford et al., 2021) and VideoCLIP (Wang et al., 2023a) text similarity scores. This shows the contribution of each component in achieving high-quality 3D generation.
Figure 5: Generation Results With and Without View Selection. The top four rows show 3D generation results for a colorful rainbow fish. The first row shows the 3D assets generated by Flex3D (displayed as rendered views) without our selection pipeline. The blue circles highlight regions exhibiting generation failures. The second row presents the generated views after applying view selection. The third and fourth rows show the sampled input views, where a green checkmark indicates that our method selected the view, and a red cross indicates that the view was rejected. The bottom four rows show the same process for a different object.
@@
Figure 5 demonstrates the impact of view selection on the quality of generated 3D assets. When view selection is not used, some of the generated input views in stage 1 may be less desirable. The blue circles highlight regions where deficiencies in the input result in poor generation quality. However, by incorporating view selection, the model is able to select the most high-quality and consistent views as input, generally resulting in improved 3D asset generation.
@@
We analyze our imperfect data simulation strategy using both reconstruction and generation metrics. Evaluation settings mirror those used in the ablation study. As shown in table 5, incorporating imperfect data simulation yields improvements across both generative and reconstruction tasks. In contrast, simpler augmentation techniques like adding Gaussian or Salt-and-pepper noise with different intensities are found to degrade reconstruction performance. This suggests that our strategy effectively exposes the model to a wider range of data variations, enhancing its overall performance and robustness.
@@ 6 Conclusion
This paper introduces Flex3D, a novel feed-forward 3D generation pipeline that produces high-quality 3D Gaussian representations from text or single-image inputs. We propose a series of approaches to overcome the limitations of previous two-stage methods, significantly improving final 3D quality. Extensive evaluations on benchmark tasks demonstrate that Flex3D achieves state-of-the-art performance in both 3D reconstruction and generation. These results highlight the effectiveness of our approach in addressing the challenges of feed-forward 3D generation, paving the way for more robust, versatile, and accessible 3D content creation workflows for a multitude of applications.
Impact Statement
@@
Appendix A Implications for future research
Feed-forward Two-stage 3D generation.
The key insight is that we introduced a series of methods to handle imperfect multi-view synthesis results in the common two-stage 3D generation pipeline. Our whole Flex3D pipeline introduces little computational cost but yields significant performance and robustness gains, and it could serve as a common design pipeline for future research in 3D generation. Additionally, all individual components proposed in this work can be easily adopted by future research in 3D generation to improve performance. Similarly, design ideas analogous to the Flex3D pipeline could be readily adopted for large 3D scene generation.
Feed-forward Two-stage 4D generation.
Moreover, our work could be beneficial for 4D generation, which is an even more challenging task that faces similar limitations to two-stage 3D generation pipelines. Our pipeline could be directly extended to handle 4D object generation tasks. One could first generate 64 views (16 time dimensions $\times$ 4 multi-views) by fine-tuning video-based diffusion models, then slightly modify the view selection pipeline to keep only those views consistent across multiple views and time dimensions. Then, extend FlexRM from a tri-plane to a hex-plane or additionally learn time offsets to enable 4D representation. This should yield a strong method for 4D asset generation.
Generative 3D/4D reconstruction.
Methods like Im-3D, Cat3D, and Cat4D rely on multi-view diffusion models to synthesize a large number of possible views, which are then used to fit a 3D/4D representation. Our work could inspire advancements in this area in two key ways:
View Selection: Our view selection pipeline could directly enhance these methods by filtering out inconsistent synthesized views before fitting a 3D representation, potentially leading to performance improvements. This approach can also be naturally extended to handle synthesized 4D views.
Efficiency: Fitting a 3D representation from scratch is time-consuming. The FlexRM reconstruction model we developed can process up to 32 views and has the potential to scale further. Synthesized views could first be processed through this model, which operates in under a second, to generate a strong initial 3D representation. This approach has the potential to dramatically reduce the time required for fitting a 3D representation—from the usual half an hour to under a minute. This idea could also be extended to 4D, as adapting the current 3D reconstruction pipeline to 4D is straightforward.
Leveraging 3D understanding for generation.
Keypoint matching techniques are used in this work to effectively mitigate multi-view inconsistencies. We hope this will also inspire the 3D generation community to incorporate advanced techniques from the rapidly evolving field of 3D understanding. Recent advances in deep learning have led to significant developments in matching, tracking, deep structure from motion, and scene reconstruction. These advancements offer the 3D generation community useful tools (such as pose estimation and keypoint matching), pseudo-supervision signals (e.g.,pseudo-depth supervision), and new model design ideas.
Appendix B Implementation details
@@ Appendix D Qualitative Results on 3D Reconstruction
Appendix E Failure cases
We present some failure cases of our candidate view generation and selection pipeline in fig. 9. A notable failure occurs when the input image contains floaters or small transparent objects. This leads to incorrect generation results, especially at different elevation angles. The view selection pipeline only partially removes these incorrect results. Additionally, our view selection pipeline can sometimes produce incorrect results, particularly with objects containing thin geometries. By enabling the use of an arbitrary number of high-quality views and enhancing robustness to input imperfections, Flex3D not only significantly improves the final 3D quality but also offers greater versatility. These thin components can contribute less to feature matching, making them more susceptible to incorrect selection.
Figure 9: Failure Cases. A green check mark indicates that our method selected the view, while a red cross indicates that the view was rejected. Question marks indicate incorrect selection results. The top two rows show results for a mushroom, highlighting difficulties with floaters or small transparent objects inside the input image. The bottom two rows illustrates the challenges in filtering generated views of objects with thin thin geometries.